TYPE AN EQUATION USING SLOPE-INTERCEPT FORM OR USING THE GIVEN POINT IN POINT-SLOPE FORM

(7,-6); 7X-3Y=5

Answer by Alan3354(69443)

TYPE AN EQUATION USING SLOPE-INTERCEPT FORM OR USING THE GIVEN POINT IN POINT-SLOPE FORM

(7,-6); 7X-3Y=5
--------------
y+6 = m*(x-7)
================
What's the equation for?